OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 3x + 5y = 3 x + 2y = 0

OpenStudy (anonymous):

(1,3/5)

OpenStudy (anonymous):

for the first one

OpenStudy (anonymous):

These are my answer choices. (1, 1) (6, –3) (6, 1) (1, –3)

OpenStudy (anonymous):

ya i just redid that lol

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

(1,-3)

OpenStudy (texaschic101):

x + 2y = 0 x = -2y now sub -2y in for x in the other equation 3x + 5y = 3 3(-2y) + 5y = 3 -6y + 5y = 3 -y = 3 y = -3 x = -2y x = -2(-3) x = 6 x = -6, y = -3 or (6,-3)

OpenStudy (texaschic101):

oops...it is not-6, it is 6

OpenStudy (anonymous):

dang i did something wrong then im sorry

OpenStudy (texaschic101):

if I had $1 for every mistake I made...I would be rich...lol

OpenStudy (anonymous):

Its okay I did it and I didn't get the same answer. Then I got the same answer as @texaschic101 .